40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 19 23:44:00 2018, gautam, Update, IOO, IMC loop checkup IMC_RF_noise.pdfIMC_OLTF_20180320.pdf
    Reply  Tue Mar 20 16:53:03 2018, gautam, Update, IOO, IMC loop checkup IMC_RF_noise.pdf
       Reply  Tue Mar 20 21:08:03 2018, gautam, Update, IOO, IMC loop checkup 
          Reply  Tue Mar 20 22:44:45 2018, gautam, Update, IOO, IMC loop checkup 
             Reply  Wed Mar 21 15:52:45 2018, gautam, Update, IOO, MC error point calibration IMC_PDHdisc_20180321.pdfIMC_PDHdisc785_20180321.pdfIMC_oscope.pdf
                Reply  Wed Mar 21 17:31:21 2018, gautam, Update, IOO, MC error point calibration IMC_RF_noise_calib.pdf
                   Reply  Wed Mar 21 21:13:44 2018, gautam, Update, IOO, IMC noise budget IMC_RF_noise_calib.pdf
                      Reply  Tue Mar 27 01:39:44 2018, gautam, Update, IOO, PSL noise eater was off IR_ALS_20180326.pdf
                         Reply  Tue Mar 27 19:32:03 2018, arijit, Update, IOO, PSL noise eater was off MCREFL_TF.pdfMCREFL_SPECTRUM.pdf
                            Reply  Tue Mar 27 23:37:35 2018, gautam, Update, IOO, MC REFL PD removed IMG_6955.JPG
                               Reply  Wed Mar 28 21:31:39 2018, gautam, Update, IOO, MC REFL PD removed 
                            Reply  Wed Mar 28 21:47:37 2018, arijit, Update, IOO, MCREFL_PD Optical response measurement LISO_fit_with_info.pdfLISO_fit.zip
                               Reply  Thu Mar 29 17:57:36 2018, arijit, Update, IOO, MCREFL_PD Optical response measurement NOISE_EATER_On_OFF.pdf
                                  Reply  Fri Mar 30 06:14:31 2018, rana, Update, IOO, MCREFL_PD Optical response measurement 
                                     Reply  Fri Mar 30 22:37:36 2018, Kevin, Update, IOO, MCREFL_PD Optical response measurement REFLPD_DemodBoard.pdf
                      Reply  Thu Apr 5 04:36:56 2018, Kevin, Update, IOO, Coil driver noise MC3.pdfMC2.pdfCoilDriver.pdf
                         Reply  Thu Apr 5 12:13:18 2018, Koji, Update, IOO, Coil driver noise 
                            Reply  Fri Apr 6 21:39:09 2018, gautam, Update, IOO, MC2 suspension health checkup MC2_Freeswinging.pdf
                               Reply  Mon Apr 9 18:46:03 2018, gautam, Update, IOO, Further debugging MC_Freeswinging.pdfIMC_NB_20180409.pdf
                                  Reply  Tue Apr 10 14:28:44 2018, gautam, Update, IOO, Further debugging 
                                     Reply  Wed Apr 11 01:34:31 2018, gautam, Update, IOO, Activities today 
                                        Reply  Thu Apr 12 18:12:49 2018, gautam, Update, ALS, NPRO channels hijacked IR_ALS_20180412.pdfBeatMouthDrift.pngETMX_20180416.jpg
                                           Reply  Fri Apr 13 11:02:41 2018, gautam, Update, ALS, EX fiber polarization drift EX_pol_drift.png
                                              Reply  Fri Apr 13 17:56:26 2018, gautam, Update, ALS, Fibers switched out 
                                                 Reply  Sat Apr 14 14:42:09 2018, gautam, Update, ALS, Fibers switched out PSLdrift.png
                                                    Reply  Tue Apr 17 14:08:29 2018, gautam, Update, ALS, Fibers switched out PSL_fluctuations.png
                                                       Reply  Fri Apr 20 00:26:34 2018, gautam, Update, ALS, Fibers switched out 
                                                          Reply  Sat Apr 21 20:25:12 2018, gautam, Update, ALS, PSL fiber pickoff status PSL-beatMouthPickoff.png
                                  Reply  Thu Apr 19 00:03:51 2018, gautam, Update, IOO, More IMC NBing IMC_NB_20180409.pdfIMC_coils_20180418.pdf
                                     Reply  Thu Apr 19 17:15:35 2018, gautam, Update, IOO, More IMC NBing IMC_NB_20180419.pdfIMC_coilDriverNoises_20180419.pdf
                                        Reply  Sun May 20 23:43:50 2018, gautam, Update, IOO, Coil driver noise re-measurement MC1_coilDriver.pdfMC2_coilDriver.pdfMC3_coilDriver.pdfMC_coilDriverNoises.tgzActuationChainNoiseMeas.pdf
                                           Reply  Tue May 22 17:26:25 2018, gautam, Update, IOO, MC1 Coil Driver pulled out 
                                              Reply  Tue May 22 23:28:01 2018, gautam, Update, IOO, MC1 Coil Driver pulled out 
                                                 Reply  Wed May 23 17:58:48 2018, gautam, Update, IOO, MC1 Coil Driver pulled out MC1_monitorTF.pngMC1_ULnoise.pdf
                                                    Reply  Wed May 30 10:17:46 2018, gautam, Update, IOO, MC1 Coil Driver pulled out EM_pickup.pdfcoilDriverNoiseComparison.pdf
                            Reply  Fri Apr 6 22:23:53 2018, Kevin, Update, IOO, Coil driver noise MC2_coil_driver.pdf
Message ID: 13773     Entry time: Fri Apr 20 00:26:34 2018     In reply to: 13757     Reply to this: 13779
Author: gautam 
Type: Update 
Category: ALS 
Subject: Fibers switched out 

Summary:

I think the dominant cause for the fact that we were seeing huge swing in the power coupled into the fiber was that the beam being sent in was in fact not linearly polarized, but elliptically polarized. I've rectified this with the help of a PBS. Fiber has been plugged into my polarization monitoring setup. Let's monitor for some long stretch and see if the situation has improved.

Details:

  • The new fiber mount I ordered, K6XS, arrived today. I like it - it has little keys with which all DoFs can be locked. Moreover, it is compatible with the fixed collimators which IMO is the easiest way to achieve good mode-matching into the fiber. It is basically a plug-and-play replacement for the mounts we were using. Anyways, we can evaluate the performance over the coming days.
  • I installed it on the PSL table (started work ~10pm, HEPA turned up to maximum, PSL shutter closed).
  • But even with the new rotational DoF locking feature, I saw that slight disturbances in the fiber caused wild fluctuations in my polarization monitoring setup PD outputs. This was a useful tool through the night of checking the polarization content in the two special axes - Aidan had suggested using a heat gun but shaking the fiber a bit works well too I think.
  • The PM980 fiber has an alignment key that is aligned with the slow axis of the fiber - so it is a useful alignment reference. But even by perturbing the roational alignment about the vertical by +/-15 degrees, I saw no improvement in this behavior. So I began to question my assumption that the input beam itself had clean polarization content.
  • Since my pickoff beam has gone through a QWP and two PBSs, I had assumed that the beam was linearly polarized.
  • But by putting a PBS just upstream of the input fiber coupler, I could see a beam at the S-port with an IR card (while I expected the beam to be P-polarized).
  • OK - so I decided to clean up the input polarization by leaving this PBS installed. With this modification to the setup, I found that me shaking the fiber around on the PSL table didn't affect the output polarization content nearly as dramatically as before!!yes
  • The state I am leaving it in tonight is such that there is ~100x the power in the P-polarization output monitor as the S-polarization (PER ~ 20dB). I didn't try and optimize this too much more for now, I want to observe some long term trend to see if the wild power fluctuations have been mitigated.
  • The output coupler is mounted on the inferior K6X mount, and so there is the possibility that some drift will be attributable to rotation of the output coupler in it's mount. Thermally driven length changes / time varying stresses in the fiber may also lead to some residual power fluctuations. But I don't expect this to be anywhere near the ~25% I reported in the previous elog.
  • The rejected beam from the PBS was measured to be ~300 uW. I haven't dumped this properly, to be done tomorrow.
  • HEPA turned back down to 30%, PSL enclosure closed up, PSL shutter re-opened ~0030am.
  • Note that the EX and EY fiber coupled beams are also likely subject to the same problem. We have to double check. I think it's better to have a PBS in front of the input fiber coupler as this also gives us control over the amount of light coupled into the fiber.

Power budget:

Power in Measured power (Ophir, filter OFF)
@Input coupler, before PBS 4.4 mW
P-pol content @ input coupler 4.06 mW
S-pol (rejected) from PBS 275 uW
@Output coupler 2.6 mW (MM ~65%)

 

ELOG V3.1.3-